Falconmaster:

What is the fifth term of a sequence whose first term is 5 and whose common ratio is 3? 243 405 1,215

EndersWorld:

405

aestheticmorty:

B or 405

EndersWorld:

5*3=15*3=45*3=135*3=405

Falconmaster:

i owe you both my soul

EndersWorld:

I already have it ( ͡° ͜ʖ ͡°)

Falconmaster:

its worth about a penny a paper clip and pocket lint